the measure of the supplement of an angle is 20 more than three times the measure of the original apple. find the messures of the angles.
Answer by nerdybill(7384)

the measure of the supplement of an angle is 20 more than three times the measure of the original apple. find the messures of the angles.
.
Let x = original angle
then
3x+20 = supplement
.
x + (3x+20) = 180
4x+20 = 180
4x = 160
x = 40 degrees
.
supplement:
3x+20 = 3(40)+20 = 120+20 = 140 degrees
.
answer: 40 and 140 degrees